OpenStudy (anonymous):

2x+2y=19 and 2x-2y=5 solve using the addition method

OpenStudy (jdoe0001):

add them up, what do you get? \(\begin{array}{llll} 2x+2y=19\\ 2x-2y=5\\ \hline\\ \square x+\square y=\square \end{array}\) ?